The rate of a particular reaction quadruples when the temperature is increased from 10.°C to 45°C. Calculate the activation energy for this reaction.
1
Expert's answer
2012-10-05T08:16:37-0400
K = A*e^(-Ea/RT)

K1/K2= e^(Ea*(1/RT2 - 1/RT1))

Ea=(ln (K1/K2)) / (1/RT2 - 1/RT1)

Ea=(ln(1/4))/(1/8.31*318 -1/8.31*283)=29.62 kJ